# Diffusion differentiable resampling

### v2

@@ 5 Related work
In concurrent work, Gourevitch et al. (2026) propose a differentiable reparameterisation of categorical distributions, constructed using stochastic interpolants. While our diffusion resampler similarly constitutes a differentiable relaxation of categorical sampling, the key departure is that the closed-form denoiser in their work is derived under one-hot encoded samples $\{X_{i}\}_{i=1}^{N}$, whereas we focus on samples in $\mathbb{R}^{d}$. The two approaches coincide when the samples are discrete variables and when formulated within the same diffusion model framework.
@@ 6 Conclusion
Limitations and future work. We observe that propagating gradients through the diffusion resampling can be numerically sensitive to the SDE solver. This could potentially be mitigated by, e.g., reversible adjoint Heun’s method (Kidger et al., 2021) or related variants. We have also mentioned the possibility to speed up the computation via a Gibbs sampler which warrants future development. We finally note that the ensemble score may not be the only choice for diffusion resampling, as long as the strategy in Remark 1 retains.
@@ Acknowledgements
Authors’ contributions are as follows. JA verified the theoretical results, wrote Related work, and conducted the pendulum experiment. ZZ came up with the idea and developed the method, wrote the initial draft, proved the theoretical results, and performed the other experiments. All authors jointly revised the final manuscript.
@@ Appendix C Proof of Corollary 1
Note that the factor $N^{-c}$ is likely suboptimal since it can not exceed the importance sampling factor $r$. This is intuitive, because we are essentially asking $N$ to compensate the SDE accumulation error. It is possible to relax from this and choose $c\geq r$. This in turn means that we need more assumptions on $C_{e}$ to ensure that $\overline{C}_{e}(t,T(t))$ is non-increasing in $t$, in contrast to what we aimed here for being increasing which is mild. However, such settings will make Assumption 2 more difficult to verify in reality. Since making mild and practically checkable assumptions on score approximation remains an open discussion (see, e.g., De Bortoli, 2022; Chen et al., 2023a), we opt for a suboptimal bound under assumptions that are simple and easy-to-verify (n.b., we have only assumed $C_{e}$ to be a positive non-increasing function without specifying any rate).
@@ Appendix D Elaboration of Remark 1
Appendix E Common experiment settings
@@
Appendix F Gaussian mixture resampling
@@
Appendix G Time comparison
@@
Table 5: The resampling error of diffusion resampling in terms of the number of time steps $K$ and sample size $N$. Related to the time experiment in Appendix G.
@@
Table 6: The resampling error of OT in terms of entropy regularisation $\varepsilon$ and sample size $N$. Related to the time experiment in Appendix G.
@@
Appendix H Linear Gaussian SSM
@@
Appendix I Prey-predator model
@@
We simulate the model with Milstein’s method and generate data at $t\in[0,3]$ with 256 dicretisation steps. The discretisation results in an SSM $Z_{j+1}=f(Z_{j},\epsilon_{j})$, where $Z_{j}\in\mathbb{R}^{2}$ encodes $C(t_{j})$ and $R(t_{j})$, and $\epsilon_{j}\in\mathbb{R}^{2}$ stands for Brownian motion increment. We train a neural network to approximate the SDE dynamics by minimising the negative log-likelihood, estimated by a particle filter with resampling. The neural network is simply a two-layers fully connected network with residual connection, see Figure 7, mimicking a discrete SDE solver. No prior knowledge of the dynamics is incorporated into the neural network. We train the neural network at a fixed 1,000 iterations by a Lion optimiser (Chen et al., 2023b) with learning rate 0.005. At testing, we make 100 independent predictions from the learnt model and compute the root mean square error (RMSE) with respect to a reference trajectory generated using the true dynamics. We use 64 particles, and repeat the experiments 20 times.
@@
OT ($\varepsilon=0.3$) $5.62\pm 5.12$ 6
OT ($\varepsilon=0.5$) $5.69\pm 5.31$ 6
OT ($\varepsilon=1.0$) $5.30\pm 5.07$ 7
OT ($\varepsilon=1.5$) $6.44\pm 5.39$ 5
Gumbel (0.1) NaN 0
Gumbel (0.3) $4.29\pm 3.77$ 9
Gumbel (0.5) $7.48\pm 6.86$ 5
Soft (0.5) $3.93\pm 3.28$ 4
Soft (0.7) $3.85\pm 2.76$ 4
Soft (0.9) $4.38\pm 2.78$ 4
@@
Appendix J Vision-based pendulum dynamics tracking
@@
As we have seen, even in this challenging setting, diffusion resampling remains competitive among the baselines. Figure 10 shows a qualitative comparison of a top performing model in each resampling class in this setting. Here, the diffusion resampler is among the two top-performing baselines. While soft resampling occasionally shows better performance in terms of final metrics compared to diffusion resampling, see Figure 9, we note that it also exhibits significantly worse weight degeneracy compared to diffusion resampling and the other baselines. In addition, it is by construction not fully differentiable (see Table 17).
@@
Table 16: Prediction quality (SSIM, PSNR) for the pendulum dynamics tracking experiment, with respect to models trained using OT, Gumbel and Soft resampling. Results correspond to the second experiment setting. For OT, only 1 run (*) was non-divergent.
@@
OT ($\varepsilon=0.5$) $0.697^{*}$ $17.4^{*}$
OT ($\varepsilon=1.0$)
OT ($\varepsilon=1.5$)
@@
Figure 8: (Left) Median loss during the training process for the top performing configuration in each resampling class in the first setting. We note that diffusion resampling achieves the lowest loss. (Right) Median loss during the training process for the top performing configuration in each resampling class in the more challenging setting. We note that for OT, the reported median loss is based on the single successful run and is therefore not statistically meaningful. While Soft achieves the lowest median loss (see the text for discussion), diffusion resampling remains competitive and yields low, stable loss, whereas Gumbel performs markedly worse than all baselines.
Figure 9: Mean prediction SSIM and PSNR for the best (by mean) model configuration of each resampler in the more challenging setting. Individual runs are shown as scatter points. We find that although Soft can occasionally give better results than diffusion resampling, its results exhibit large variability. We also observe that Gumbel performs worse than diffusion resampling.
Figure 10: Qualitative comparisons of the learnt pendulum dynamics trained end-to-end using a differentiable resampler in the SMC training loop. The results correspond to the second, more challenging experiment setting using resampling at nearly all filtering steps. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom). Each panel shows a top result in terms of mean SSIM for each differentiable resampling method in our comparison: a) Diffusion (mean SSIM/PSNR $0.761/17.0$, b) Soft ($0.816/19.2$), c) OT ($0.697/17.4$), d) Gumbel ($0.663/16.9$).
@@
Appendix K Additional related work
@@
In addition to the discussion in Section 5, there are also connections to other, less central, lines of work. For instance, the ensemble score estimator is related to the kernel projection used in Stein variational gradient flow (Liu, 2017). The controlled gain term in Yang et al. (2013) also plays a role analogous to resampling in a continuous-time limit, but involves a computationally demanding PDE-solving step. While there are interesting connections to our work, the approach focuses on continuous-time flow filtering (cf. Kang et al., 2025) and leads to algorithms that are no longer standard SMC.
Appendix L Take-away messages
@@
  • Diffusion resampling is useful not only for differentiation, but also for reducing resampling error in general. The primary reason for this is that diffusion resampling can take additional information (e.g., $\pi_{\mathrm{ref}}$) of the target into account. When we have a complete information $\pi_{\mathrm{ref}}=\pi$, the diffusion resampling becomes an optimal resampling algorithm. Corollary 1 explicitly shows how the resampling error can be controlled depending on how well $\pi_{\mathrm{ref}}$ approximates $\pi$, calibrated by $b$. In contrast, multinomial resampling only uses information from the given samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ which uses incomplete and less information compared to diffusion resampling based on first principle.

### v3

@@ 5 Related work
In concurrent work, Gourevitch et al. (2026) propose a differentiable reparameterisation of categorical distributions using stochastic interpolants. While our diffusion resampler similarly constitutes a differentiable relaxation of categorical sampling, the key departure is that their closed-form denoiser is derived under one-hot encoded samples $\{X_{i}\}_{i=1}^{N}$, whereas we focus on samples in $\mathbb{R}^{d}$. We further consider the case when the empirical measure approximates an underlying continuous distribution as $N\to\infty$, while Gourevitch et al. (2026) mostly focus on discrete categorical distributions. There exists no exact reparametrisation producing the true expectation gradient under the categorical measure, however, a consistent reparametrisation (e.g., ours) exists converging to the true expectation gradient under the underlying continuous measure.
@@ 6 Conclusion
Limitations and future work. We observe that propagating gradients through the diffusion resampling can be numerically sensitive to the SDE solver. This could potentially be mitigated by, e.g., reversible adjoint Heun’s method (Kidger et al., 2021) or related variants. We also note that the ensemble score may not be the only choice to achieve diffusion resampling.
@@ Acknowledgements
Authors’ contributions are as follows. JA verified the theoretical results, wrote Related work, and conducted the pendulum and weather forecasting experiments. ZZ came up with the idea and developed the method, wrote the initial draft, proved the theoretical results, and performed the other experiments. All authors contributed to and revised the final manuscript.
@@ Appendix C Proof of Corollary 1
Note that the factor $N^{-c}$ is likely suboptimal since it can not exceed the importance sampling factor $r$. This is intuitive, because we are essentially asking $N$ to compensate the SDE accumulation error. It is possible to relax from this and choose $c\geq r$. This in turn means that we need more assumptions on $C_{e}$ to ensure that $\overline{C}_{e}(t,T(t))$ is non-increasing in $t$, in contrast to what we aimed here for being increasing which is mild. However, such settings will make Assumption 2 more difficult to verify in reality. Since making mild and practically checkable assumptions on score approximation remains an open discussion (see, e.g., De Bortoli, 2022; Chen et al., 2023), we opt for a suboptimal bound under assumptions that are simple and easy-to-verify (n.b., we have only assumed $C_{e}$ to be a positive non-increasing function without specifying any rate).
@@ Appendix D Elaboration of Remark 1
Appendix E Error analysis of the resampling mapping
Previously in our main results (e.g., Corollary 1), we have analysed the error between the resampled distribution and the underlying true continuous distribution. In this section, we forgo the continuous distribution, and purely analyse the resampling mapping, that is, the $L_{2}$ error between the input and output ensembles.
Recall the input ensemble $\pi^{N}\coloneqq\sum_{i=1}^{N}w_{i}\,\delta_{X_{i}}$, the resampled ensemble $\pi^{*,N}$, and denote $\pi^{N}(\phi)\coloneqq\mathbb{E}_{\pi^{N}}[\phi]=\sum_{i=1}^{N}w_{i}\,\phi(X_{i})$ for any test function $\phi$. Define the $\chi^{2}$ divergence by $\chi^{2}(p\,\|\,q)\coloneqq\int(p(x)-q(x)\,/\,q(x)^{2})\,q(x)\mathop{}\!\mathrm{d}x$. Based on Equation (30) we have
$\pi^{\star,N}(\phi)=\int\pi_{\mathrm{ref}}(u_{0})\sum_{i=1}^{N}\phi(X_{i})\,\frac{w_{i}\,p_{T|0}(u_{0}{\;|\;}X_{i})}{p_{T}^{N}(u_{0})}\mathop{}\!\mathrm{d}u_{0}=\int\pi_{\mathrm{ref}}(u_{0})\,\tilde{\phi}_{N}(u_{0})\mathop{}\!\mathrm{d}u_{0},$ (31)
where $\tilde{\phi}_{N}(u_{0})\coloneqq\sum_{i=1}^{N}\phi(X_{i})\,\frac{w_{i}\,p_{T|0}(u_{0}{\;|\;}X_{i})}{p_{T}^{N}(u_{0})}$. Then $\pi^{N}(\phi)=\sum_{i=1}^{N}w_{i}\,\phi(X_{i})\int p_{T|0}(u_{0}{\;|\;}X_{i})\mathop{}\!\mathrm{d}u_{0}=\int p_{T}^{N}(u_{0})\,\tilde{\phi}_{N}(u_{0})\mathop{}\!\mathrm{d}u_{0}$ and we have the residual $\pi^{\star,N}(\phi)-\pi^{N}(\phi)=\int(\pi_{\mathrm{ref}}(u_{0})-p_{T}^{N}(u_{0}))\,\tilde{\phi}_{N}(u_{0})\mathop{}\!\mathrm{d}u_{0}$. Hence
$\bigl\lvert\pi^{\star,N}(\phi)-\pi^{N}(\phi)\bigr\rvert^{2}=\int\biggl\lvert\tilde{\phi}_{N}(u_{0})\,\bigl(\pi_{\mathrm{ref}}(u_{0})-p_{T}^{N}(u_{0})\bigr)\,\frac{\sqrt{p_{T}^{N}(u_{0})}}{\sqrt{p_{T}^{N}(u_{0})}}\biggr\rvert\mathop{}\!\mathrm{d}u_{0}\leq\pi^{N}(\phi^{2})\,\chi^{2}(\pi_{\mathrm{ref}}\,\|\,p_{T}^{N}),$ (32)
where the identity $\tilde{\phi}_{N}(u_{0})^{2}\leq\sum_{i=1}^{N}\phi(X_{i})^{2}\,\frac{w_{i}\,p_{T|0}(u_{0}{\;|\;}X_{i})}{p_{T}^{N}(u_{0})}$ was applied. Assume that the test function $\phi$ is uniformly bounded, i.e., $\pi^{N}(\phi^{2})\leq c_{\phi}$, and also knowing that $p_{T}=\operatorname{\mathbb{E}}[p_{T}^{N}]$, we finally have
$\operatorname{\mathbb{E}}\bigl[\lvert\pi^{\star,N}(\phi)-\pi^{N}(\phi)\rvert^{2}\bigr]\leq c_{\phi}\,\operatorname{\mathbb{E}}[\chi^{2}(p_{T}\,\|\,p_{T}^{N})]+c_{\phi}\,\operatorname{\mathbb{E}}\biggl[\int\frac{p_{T}(u)-\pi_{\mathrm{ref}}(u)}{p_{T}^{N}(u)}\mathop{}\!\mathrm{d}u\biggr].$ (33)
Appendix F Common experiment settings
@@
Appendix G Gaussian mixture resampling
@@
Appendix H Time comparison
@@
Table 5: The resampling error of diffusion resampling in terms of the number of time steps $K$ and sample size $N$. Related to the time experiment in Appendix H.
@@
Table 6: The resampling error of OT in terms of entropy regularisation $\varepsilon$ and sample size $N$. Related to the time experiment in Appendix H.
@@
Appendix I Linear Gaussian SSM
@@
Appendix J Prey-predator model
@@
We simulate the model with Milstein’s method and generate data at $t\in[0,3]$ with 256 dicretisation steps. The discretisation results in an SSM $Z_{j+1}=f(Z_{j},\epsilon_{j})$, where $Z_{j}\in\mathbb{R}^{2}$ encodes $C(t_{j})$ and $R(t_{j})$, and $\epsilon_{j}\in\mathbb{R}^{2}$ stands for Brownian motion increment. We train a neural network to approximate the SDE dynamics by minimising the negative log-likelihood, estimated by a particle filter with resampling. The neural network is simply a three-layers fully connected network with residual connection, see Figure 7, mimicking a discrete SDE solver. No prior knowledge of the dynamics is incorporated into the neural network. We train the neural network at a fixed 1,000 iterations by the Adam optimiser with learning rate 0.005. At testing, we make 100 independent predictions from the learnt model and compute the root mean square error (RMSE) with respect to a reference trajectory generated using the true dynamics. We use 64 particles, and repeat the experiments 20 times.
@@
OT ($\varepsilon=0.3$) $2.35\pm 0.95$ 19
OT ($\varepsilon=0.5$) $2.72\pm 1.60$ 20
OT ($\varepsilon=1.0$) $3.27\pm 2.71$ 20
OT ($\varepsilon=1.5$) $3.52\pm 3.65$ 20
Gumbel (0.1) $9.05\pm 8.65$ 15
Gumbel (0.3) $2.09\pm 1.13$ 20
Gumbel (0.5) $2.40\pm 1.05$ 20
Soft (0.5) $3.29\pm 4.83$ 17
Soft (0.7) $2.13\pm 0.87$ 17
Soft (0.9) $1.89\pm 0.64$ 16
Stopped $1.96\pm 0.55$ 20
@@
Appendix K Vision-based pendulum dynamics tracking
@@
As we have seen, even in this challenging setting, diffusion resampling remains competitive among the baselines. Figure 10 shows a qualitative comparison of a top performing model in each resampling class in this setting. Here, the diffusion resampler is among the three top-performing baselines. While soft resampling occasionally shows better performance in terms of final metrics compared to diffusion resampling, see Figure 9, we note that it also exhibits significantly worse weight degeneracy compared to diffusion resampling and the other baselines. In addition, it is by construction not fully differentiable (see Table 20). We also note that while OT produces the top performing model here, it exhibits high variability and is overall unstable in this setting, with only a few converging training runs.
@@
Table 16: Prediction quality (SSIM, PSNR) for the pendulum dynamics tracking experiment, with respect to models trained using OT, Gumbel and Soft resampling. Results correspond to the second experiment setting. For OT configurations, only 3 (**) and 4(*) runs were non-divergent.
@@
OT ($\varepsilon=0.5$) $0.662^{*}\pm 0.141$ $17.7^{*}\pm 1.782$
OT ($\varepsilon=1.0$) $0.575^{**}\pm 0.303$ $17.5^{**}\pm 1.924$
OT ($\varepsilon=1.5$) $0.708^{**}\pm 0.159$ $18.2^{**}\pm 2.73$
@@
Figure 8: (Left) Median loss during the training process for the top performing configuration in each resampling class in the first setting. We note that diffusion resampling achieves the lowest loss. (Right) Median loss during the training process for the top performing configuration in each resampling class in the more challenging setting. We note that for OT, the reported median loss is based on the few successful runs and is therefore not statistically meaningful. While Soft achieves the lowest median loss (see the text for discussion), diffusion resampling remains competitive and yields low, stable loss, whereas Gumbel performs markedly worse than all baselines.
Figure 9: Mean prediction SSIM and PSNR for the best (by mean) model configuration of each resampler in the more challenging setting. Individual runs are shown as scatter points. We find that although Soft can occasionally give better results than diffusion resampling, its results exhibit large variability. We also observe that Gumbel performs worse than diffusion resampling. We further observe that OT gets top SSIM and PSNR here, but note that it exhibits large variability and is not stable in this task (most OT runs diverged during training).
Figure 10: Qualitative comparisons of the learnt pendulum dynamics trained end-to-end using a differentiable resampler in the SMC training loop. The results correspond to the second, more challenging experiment setting using resampling at nearly all filtering steps. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom). Each panel shows a top result in terms of mean SSIM for each differentiable resampling method in our comparison: a) Diffusion (mean SSIM/PSNR $0.761/17.0$, b) Soft ($0.816/19.2$), c) OT ($0.820/20.1$), d) Gumbel ($0.663/16.9$).
@@
Appendix L Bayesian neural network training
We apply SMC for a larger-scale problem: training a (partial) Bayesian neural network (Zhao et al., 2024; Sharma et al., 2023, pBNN). There are two steps moving from training a classical BNN to training a pBNN with SMC sampler. First, pBNNs only have priors on a (small) subset of the neural network parameters. This results in a latent-variable model to learn from the data, where computing the posterior distribution is easier than that of the full BNN. Let us denote the pBNN by $x\mapsto f_{\theta,z}(x)$, where $\theta$ and $z$ stand for deterministic and stochastic parameters, respectively. Second, the prior construction is no longer static or explicit (e.g., $z$ being a Gaussian). Instead, $z$ is assumed to follow a Markovian dynamics motivated by Zhao et al. (2024); Chang et al. (2022); Freitas et al. (2000) so that data are modelled as independent conditionally on a latent process. Specifically, we apply the pBNN setting and construct an SSM
$\begin{split}p(z_{j}{\;|\;}z_{j-1})&=\mathrm{N}(z_{j};\rho\,z_{j-1},1-\rho^{2}),\\ p_{\theta}(D_{j}{\;|\;}z_{j})&=\mathrm{Softmax}(f_{\theta,z_{j}};D_{j}),\end{split}$ (38)
for CIFAR10 classification, where $D_{j}$ is a batch data at surrogate time $j$ (training step), and we choose $\rho=0.99$.
We choose a ResNet18 neural network, and set the first two convolution layers be stochastic (i.e., the neural network parameters in this part is the latent variable $z$). This results in latent dimension $d_{z}=1,856$ and parameter dimension $d_{\theta}=11,172,106$. We choose the number of particles to be 8, resampling threshold 0.5, and apply an Adam optimiser train for 200 epochs. The diffusion resampling parameter is $T=1$ and $K=4$ using the Jentzen–Kloeden SDE integrator. The results of the classification evaluated using accuracy and F1 score is shown in Table 17.
Table 17: CIFAR10 classification with different resampling methods. Results are averaged over 5 independent trainings.
Diffusion OT ($\varepsilon=1$) Gumbel (0.3) Soft (0.7) Stopped
Accuracy $\mathbf{87.81}\pm 0.64$ $83.28\pm 0.63$ $86.92\pm 0.58$ $86.57\pm 0.58$ $86.28\pm 0.57$
F1 score $\mathbf{87.76}\pm 0.63$ $83.18\pm 0.70$ $86.89\pm 0.57$ $86.54\pm 0.57$ $86.21\pm 0.56$
Appendix M Weather forecast
Experiment settings
As another large-scale experiment, we consider a high-dimensional weather forecasting task to further demonstrate that the diffusion resampler scales to high-dimensional state space models. Given partial noisy observations, we track the evolution of the 850 hPa atmospheric temperature field over time. We use SMC to enable learning the parameters $\theta$ of an SSM with atmosphere dynamics model $p_{\theta}(z_{k}{\;|\;}z_{k-1})$ and observation model $p(y_{k}{\;|\;}z_{k})$. Here, the latent state $z_{k}\in\mathbb{R}^{2048}$ represents the global temperature field at $5.625^{\circ}$ resolution, with the resulting latent dimension $d=2048$ ($32\times 64$ spatial grid). The partial observation $y_{k}$ consists of 8080% of pixels selected uniformly at random at each time step, corrupted by Gaussian noise with standard deviation $\sigma_{y}=0.01$. The observation model is $p(y_{k}{\;|\;}z_{k})=\mathcal{N}(y_{k}{\;|\;}\mathcal{M}_{k}z_{k},\sigma_{y}^{2}I)$, where $\mathcal{M}_{k}$ denotes the random mask operator for the observed pixels at time step $k$. A UNet parametrises the transition density $p_{\theta}(z_{k}{\;|\;}z_{k-1})$, and is learnt using the negative log-likelihood estimate provided by a particle filter with resampling at every step.
The models are trained using ERA5 data from the WeatherBench benchmark dataset (Rasp et al., 2020). We present results for two different settings: a full year setting, where the model is tasked with iterative 4-day forecasts over a full year, and a single-sequence setting, where the model is tasked with iterative 8-day forecasts in January.
Training details
We use data from the years 2010-2018 for training (excluding 2016), with 2014 held out for evaluation. The temperature fields are normalised to $[0,1]$ using global min-max normalisation computed across all years. The UNet takes the current state $z_{k-1}$ and process noise concatenated along the channel dimension. The output is given by $z_{k}=\mathrm{Sigmoid}\bigl(z_{k-1}+\delta_{\theta}(z_{k-1},q_{k})\bigr)$, where $\delta_{\theta}$ is the UNet output, and $q_{k}\sim\mathcal{N}(0,\sigma_{q}^{2}I)$ with $\sigma_{q}^{2}=0.1$. We train the model for 3000 iterations using the Adam optimiser with learning rate $2\times 10^{-4}$ and gradient clipping with global norm threshold $1.0$. The optimiser omits at most 10 consecutive non-finite parameter updates by leaving the current parameter values unchanged. We use $N=16$ particles and trigger resampling at each filtering step. By default, the diffusion resampler jitters the diagonal of the Gaussian reference by $10^{-5}$ for numerical stability, and the log-likelihood at each time step is normalised by the number of observed pixels.
Results and evaluation
The learnt dynamics models are evaluated on unobserved (masked) pixels from the training sequences and on full sequences from the held-out year. We use the mean square error (MSE) averaged over time steps and 16 individual rollout samples to evaluate the performance. All MSE values are reported at scale $\times 10^{-3}$, averaged over 5 independent training runs. Table 18 shows results for the full year setting, where training sequences of length 16 are randomly sampled from the full year of 6-hourly data across all 7 training years. The model is evaluated on 5 evenly spaced windows from the held-out year, covering all seasons. Table 19 shows results for the single-sequence setting, where the training data consists of fixed sequences of length 32 covering the first 8 days of January for each year in the training data. The model is evaluated on the corresponding January sequence from the held-out year.
Table 18: Prediction quality in terms of MSE ($\times 10^{-3}$, lower is better) for the weather forecasting experiment in the full year setting. Results are presented as mean $\pm$ standard deviation over 5 independent runs recorded after 3000 training iterations. Train MSE is averaged over 5 evenly spaced windows from each of the 7 training years; eval MSE is computed on 5 evenly spaced windows from the held-out year.
Method Train MSE (masked) Eval MSE (full)
Diffusion, Euler–Maruyama ($T=1,K=4$,ODE) $2.2\pm 0.7$ $2.1\pm 0.8$
Diffusion, Euler–Maruyama ($T=1,K=4$,SDE) $2.5\pm 0.8$ $2.4\pm 0.7$
Diffusion, Euler–Maruyama ($T=1,K=8$,ODE) $2.1\pm 0.5$ $2.0\pm 0.4$
Diffusion, Jentzen–Kloeden ($T=1,K=8$,ODE) $2.0\pm 0.5$ $1.9\pm 0.4$
Soft ($0.7$) $1.7\pm 0.1$ $1.7\pm 0.1$
Soft ($0.9$) $1.6\pm 0.1$ $1.7\pm 0.2$
Gumbel ($0.1$) $4.6\pm 2.0$ $4.7\pm 2.3$
Gumbel ($0.3$) $1.7\pm 0.1$ $1.7\pm 0.1$
OT ($\varepsilon=0.5$) $1.9\pm 0.1$ $1.8\pm 0.1$
OT ($\varepsilon=1.0$) $1.8\pm 0.2$ $1.8\pm 0.1$
OT ($\varepsilon=1.5$) $2.4\pm 1.2$ $2.4\pm 1.4$
Table 19: Prediction quality in terms of MSE ($\times 10^{-3}$, lower is better) for the weather forecasting experiment in the single-sequence setting. Results are presented as mean $\pm$ standard deviation over 5 independent runs recorded after 3000 training iterations. Train MSE is averaged over all 7 training years; eval MSE is computed on the January sequence from the held-out year.
Method Train MSE (masked) Eval MSE (full)
Diffusion, Euler–Maruyama ($T=1,K=4$,ODE) $0.6\pm 0.1$ $1.9\pm 0.1$
Diffusion, Euler–Maruyama ($T=1,K=4$,SDE) $0.6\pm 0.1$ $2.0\pm 0.1$
Diffusion, Euler–Maruyama ($T=1,K=8$,ODE) $0.5\pm 0.0$ $2.0\pm 0.0$
Diffusion, Jentzen–Kloeden ($T=1,K=8$,ODE) $0.5\pm 0.0$ $2.0\pm 0.1$
Soft ($0.7$) $0.5\pm 0.1$ $2.1\pm 0.1$
Soft ($0.9$) $0.5\pm 0.1$ $2.0\pm 0.1$
Gumbel ($0.1$) $0.6\pm 0.1$ $2.0\pm 0.0$
Gumbel ($0.3$) $0.6\pm 0.1$ $1.9\pm 0.0$
OT ($\varepsilon=0.5$) $0.6\pm 0.1$ $1.9\pm 0.1$
OT ($\varepsilon=1.0$) $0.6\pm 0.0$ $1.9\pm 0.1$
OT ($\varepsilon=1.5$) $0.6\pm 0.1$ $2.0\pm 0.1$
Our results show that diffusion resampling can be used inside the SMC learning pipeline to learn atmosphere dynamics in this complex, high-dimensional setting. Tables 18 and Table 19 show that almost all configurations achieve relatively low MSE on both unobserved pixels in sequences drawn from the training data, and on unseen sequences drawn from the held-out evaluation data. The worst results are obtained by Gumbel ($\tau=0.1$) in the full year setting. In this setting, which contains data from the annual temperature variations, all resamplers struggle to learn fine-grained details in the current setup. Soft resampling achieves the lowest eval MSE with low variance, while the best diffusion configuration (Jentzen–Kloeden, $K=8$, ODE) achieves competitive performance. We note that these results rely on only 5 independent training runs for each model configuration. The standard deviation is not negligible in the comparison, and some configurations in the full year setting (including some diffusion configurations) exhibit relatively high variance.
Figure 12 shows examples of predicted forecasts using a model trained with diffusion resampling. These results correspond to the second experiment setting, where models are trained on a small dataset from a limited season (early January). Though Table 19 may indicate some overfitting, the evaluation forecasts show that some generalisable dynamics have been learned, and that the forecasts include fine-scale features.
Figure 12: Top panel: Training sequence forecast with corresponding masked observations (middle row). Bottom panel: Evaluation sequence forecast from the held-out year. Each panel shows ground truth (top row) and model prediction (bottom row) at evenly spaced time steps. The model was trained using the diffusion resampler in the learning pipeline.
For a more comprehensive comparison, there are limitations to address, such as the limited number of independent runs, choice of evaluation metric and the amount of data used for training. A simple MSE against the ground truth does not fully capture the predictive distribution quality and may not fully capture the forecasting performance in terms of fine-grained details and long forecasting windows. In both settings, the diffusion resampler achieves a performance comparable to the baselines. Training is stable across seeds, which shows that the diffusion resampler scales at latent dimension $d=2048$.
Appendix N Choosing the hyperparameters
Although the diffusion resampling is powerful, it comes with a variety of hyperparameters to tune: the diffusion coefficients $b$ and $\pi_{\mathrm{ref}}$, the diffusion time $T$, the integrator, and integration step $K$. Ideally, $\pi_{\mathrm{ref}}$ should well approximate the underlying continuous distribution, and the choice $b$ should correspondingly reflect the approximation error as in Corollary 1. The setting of diffusion time $T$ is arbitrary if one can sample $p_{T}$ exactly, but smaller $T$ leads to potentially lower integration steps required. As for the integrator and integration steps, our empirical results indicate that Jentzen–Kloeden and Euler–Maruyama are often a good start, and $K\leq 8$ is often sufficient for learning large neural network-parametrised systems. To retain a good computational complexity, it is suggested that the number of integration steps should be smaller than the number of samples: $K<N$.
Appendix O Additional related work
@@
In addition to the discussion in Section 5, there are also connections to other, less central, lines of work. For instance, the ensemble score estimator is related to the kernel projection used in Stein variational gradient flow (Liu, 2017). The controlled gain term in Yang et al. (2013) also plays a role analogous to resampling in a continuous-time limit, but involves a computationally demanding PDE-solving step. While there are interesting connections to our work, the approach focuses on continuous-time flow filtering (cf. Kang et al., 2025) and leads to algorithms that are no longer standard SMC. Bao et al. (2024) propose an ensemble filter where the update step was replaced by a conditional version of the diffusion in Equation (13). This too targets at Equation (30) but additionally contains errors from likelihood score approximation.
Appendix P Take-away messages
@@
  • Diffusion resampling is useful not only for differentiation, but also for reducing resampling error in general. The primary reason for this is that diffusion resampling can take additional information (e.g., $\pi_{\mathrm{ref}}$) of the target into account. When we have a complete information $\pi_{\mathrm{ref}}=\pi$, the diffusion resampling becomes an optimal resampling algorithm. Corollary 1 explicitly shows how the resampling error can be controlled depending on how well $\pi_{\mathrm{ref}}$ approximates $\pi$, calibrated by $b$. In contrast, multinomial resampling only uses information from the given samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ which contains incomplete and less information compared to diffusion resampling using the continuous distribution approximation.